OpenStudy (anonymous):

Given sin(x)=-3/5 and x is in quadrant 3, what is the value of tan(x/2)?

OpenStudy (anonymous):

draw a triangle to find the third side, or remember the mother of all right triangles 3- 4 - 5

OpenStudy (anonymous):



OpenStudy (anonymous):

then use the half angle formula to find \[\tan(\frac{x}{2})\]

OpenStudy (anonymous):

how do you use the half angle formula

OpenStudy (anonymous):

idk look it up

OpenStudy (bradely):

tan(x/2) = sinx/(1+cosx) sinx = -3/5 and cos -4/5 Substitute find tan x/2 Source: mathskey.com

OpenStudy (anonymous):

so is the answer -3

OpenStudy (bradely):

correct

OpenStudy (anonymous):

thx for the help!!!